Supplemental Report – June 4th

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and installation is pending.

Power system:
Solar: Nominal. All input breakers of MPTT were off on June 2nd. After resetting the breaker work is normal.SOC Last 24 hours: Max 100%; Min 24 %; Avg 55%.
VDC Last 24 hours: Max 58.36V; Min 42.13V; Avg 50.02V.Generator run time: used during URC for 17.4 hours total. Generator hours 3827.7.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 56%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 68%

Water (Static Tank) – 480 gallons
Water in GreenHab – 152 gallons
Water (Outpost tank) – 250 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: Yes.
Hours: 193.9
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for supplies to pick up.

Crew Car used and why, where? Yes. To/From Grand Junction by URC/crew.

General notes and comments: The crew car needs oil change. Should be done during the next trip to Grand Junction.

Dual Split in Science Dome: A/C is on for 74F.

Summary of Internet: Nominal

Summary of suits and radios: All unplugged, visors covered, all chargers packed in the spare parts box. All radio chargers are unplugged.

EVA COMMS: Nominal.

Campus wide inspection, if action taken, what and why: all cleaned after URC.

Summary of General Operations: All nominal.

Summary of Hab operations: Cleaned after URC.

Summary of GreenHab Operations: Nothing to report.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and installation is pending.

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations: Nominal.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site